# Modelling the 5G Energy Consumption using Real-world Data: Energy Fingerprint is All You Need

### v1

@@ V-D Evaluating the Effectiveness of Attention Mechanism
In this section, we assess the efficacy of our proposed ARL through a rigorous ablation study. We compared the performance of two network configurations: one incorporating the ARL and the other devoid of it. The results, as delineated in Table IV, reveal a marked improvement in MAPE across all BS when the ARL is used, at the cost of a slight increase in the model’s parameters. Importantly, integration of ARL not only reduced the MAPE for known BSs but also significantly for unknown BSs, reducing their error from 8.5095% to 6.6863%, thus highlighting its ability to improve network generalization in various scenarios.
To evaluate if such performance improvement is due to the increased number of parameters, we introduce a model variant with three hidden layers (256-128-64) that facilitate a balanced comparison. This modification aimed to establish a baseline for comparison that is fair in terms of network complexity and capability. However, this comparison shows that a mere increase in parameters does not correlate with a significant enhancement in model performance.
TABLE IV: Performance comparison of models with and without attention mechanisms on MAPE(%).
Model Parameters Unknown-BS Known-BS All-BS w/ ARL (Ours) 105209 6.6863 4.0 4.9811 w/o ARL 100097 8.5095 4.0561 5.6571 3-hidden-layer w/o ARL 159233 9.0253 4.0815 5.959
@@ VI Conclusion
In this paper, we presented an innovative deep learning model for estimating the energy consumption of 5G BSs, a critical component of the rapidly evolving 5G networks.

### v2

@@ V-D Evaluating the Effectiveness of Attention Mechanism
In this section, we assess the efficacy of our proposed ARL through an ablation study. We compare two network configurations: one with ARL and one without. The results, shown in Table II, reveal that ARL significantly improves MAPE across all BSs, with only a slight increase in model parameters. Notably, ARL reduces MAPE for unknown BSs from 8.5095% to 6.6863%, highlighting its ability to enhance generalization. To evaluate that performance gains are due to increased parameters, we compare ARL to a deeper MLP model (256-128-64). Results show that simply increasing model depth does not significantly improve performance, confirming that ARL’s gains stem from its unique design rather than model scaling.
TABLE II: Comparison of BSID encoding methods, One-Hot encoding combinations, and models with attention mechanisms on MAPE (%). Best performances are highlighted in bold.
BSID Encoding Methods Methods (Dim) Cross-domain In-domain Average Embedding BSID w/ RM(204) (Ours) 6.6863 4.0 4.9811 Embedding BSID w/o RM(204) 21.8093 3.9984 11.4294 One-Hot BSID w/ RM(1064) 9.44 3.9452 6.1313 One-Hot BSID w/o RM(1064) 20.9739 4.0732 11.2645 No BSID(140) 8.1589 9.2191 8.7538 Baseline[11](55) 13.9083 11.3856 12.7587 One-Hot Encoding Combinations Combinations (Dim) Cross-domain In-domain Average ABF (204) (Ours) 6.6863 4.0 4.9811 AB (172) 15.3823 4.5665 9.4701 AF (188) 13.0811 4.0847 7.9097 BF (199) 11.7292 4.1652 7.34 A (156) 20.094 4.7138 11.3221 B (167) 12.7202 4.2781 7.3687 F (183) 11.1229 4.3239 7.2736 Numerical (151) 13.7479 4.0733 8.3435 Models with Attention Mechanisms Model Cross-domain In-domain Average Parameters ARL (Ours) 6.6863 4.0 4.9811 105209 w/o attention 8.5095 4.0561 5.6571 100097 Multi-head Attention 9.0952 4.0944 5.6668 163073
@@ VI Conclusion
In this paper, we presented an innovative deep learning model for estimating the energy consumption of 5G BSs, designed to enhance energy-saving efforts in the rapidly evolving 5G networks.
@@
Acknowledgment
The work was supported in part by National Key R&D Program of China under Grant 2022YFA1003900, in part by Guangdong Major Project of Basic and Applied Basic Research under Grant 2023B0303000001, in part by Guangdong Young Talent Research Project (Grant No. 2023TQ07A708).